OpenStudy (anonymous):

a line passes through the point 5,-1 and has a slope of -3. write an equation in slope intercept form for this line.

OpenStudy (anonymous):

Use y1 - y = m (x1 - x), which in turn gives you y1 - (-1) = 3 (x1 - 5) and solve for y1.

OpenStudy (texaschic101):

you can use y = mx + b (5,-1) x = 5 and y = -1 slope(m) = -3 now sub all that into y = mx + b, and find b

OpenStudy (texaschic101):

or you can use the formula that derpington supplied and find it that way also...answers will be the same